# Adversarial Training on Purification (AToP): Advancing Both Robustness and Generalization

### 10 Mar 2024 v2

@@ Abstract
The deep neural networks are known to be vulnerable to well-designed adversarial attacks. The most successful defense technique based on adversarial training (AT) can achieve optimal robustness against particular attacks but cannot generalize well to unseen attacks. Another effective defense technique based on adversarial purification (AP) can enhance generalization but cannot achieve optimal robustness. Meanwhile, both methods share one common limitation on the degraded standard accuracy. To mitigate these issues, we propose a novel pipeline called Adversarial Training on Purification (AToP), which comprises two components: perturbation destruction by random transforms (RT) and purifier model fine-tuned (FT) by adversarial loss. RT is essential to avoid overlearning to known attacks resulting in the robustness generalization to unseen attacks and FT is essential for the improvement of robustness. To evaluate our method in an efficient and scalable way, we conduct extensive experiments on CIFAR-10, CIFAR-100, and ImageNette to demonstrate that our method achieves state-of-the-art results and exhibits generalization ability against unseen attacks.
@@ 1 Introduction
Deep neural networks (DNNs) have been shown to be susceptible to adversarial examples (Szegedy et al., 2014; Deng et al., 2020), which are generated by adding small, human-imperceptible perturbations to natural images, but completely change the prediction results to DNNs, leading to disastrous implications (Goodfellow et al., 2015). Since then, numerous methods have been proposed to defend against adversarial examples. Notably, adversarial training (AT) has garnered significant attention as a robust defense technique against adversarial attacks (Goodfellow et al., 2015; Madry et al., 2018a). AT has proven effective in achieving state-of-the-art robustness against known attacks, albeit with a tendency to overfit the specific adversarial examples seen during training. Consequently, within the framework of AT, the persistent conundrum revolves around striking a balance between standard accuracy and robustness (Zhang et al., 2019), a dilemma that has been a subject of enduring concern. Furthermore, the robustness of AT defense against multiple or even unforeseen attacks still remains challenging (Poursaeed et al., 2021; Laidlaw et al., 2021; Tack et al., 2022).
More recently, the concept of adversarial purification (AP) has emerged, aiming to eliminate perturbations from potentially attacked images before inputting them into the classifier model (Yang et al., 2019; Shi et al., 2021; Nie et al., 2022). In contrast to AT methods, AP operates as a pre-processing step that can defend against unseen attacks without retraining the classifier model. However, AP exhibits lower robustness against known attacks compared to AT methods. Meanwhile, both AT and AP have a common limitation on the degraded standard accuracy (Tramer et al., 2020; Chen et al., 2022), which restricts their applicability in real-world tasks. As summarized in Table 1, the present landscape of defense methods is confronted with a formidable challenge: How can we enhance robustness against known attacks while maintaining generalization to unseen attacks and preserving standard accuracy on clean examples?
@@
To tackle these challenges with the framework of AT and AP, we propose a novel defense technique that conceptionally separates AP method into two components: perturbation destruction by random transforms and fine-tuning the purifier model via supervised AT called Adversarial Training on Purification (AToP), as shown in Figure 1. Specifically, we develop three types of random transforms and conduct a comprehensive investigation into their effectiveness in augmenting robustness and generalization against attacks. The purifier model is designed to reconstruct clean examples from corrupted inputs, which can also ensure accurate classification outputs regardless of whether the inputs are corrupted adversarial examples or corrupted clean examples. Furthermore, more importantly, we introduce the adversarial loss calculated from the output of classifier model to fine-tune the purifier model in a supervised manner while keeping the classifier model fixed. Differing from the traditional AP methods, our method can learn the purifier model which can not only generate high-quality purified examples but also avoid generating examples to incorrect classes.
We empirically evaluate the performance of our method by comparing with the latest AT and AP methods across various attacks (i.e., FGSM (Goodfellow et al., 2015), PGD (Madry et al., 2018b), CW (Carlini & Wagner, 2017), AutoAttack (Croce & Hein, 2020), and StAdv (Xiao et al., 2018)) and conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette employing multiple classifier models and purifier models. The experimental results demonstrate that our method achieves state-of-the-art results and exhibits robust generalization against unseen attacks. Furthermore, our method significantly improves the performance of the purifier model in robust classification, surpassing the performance of the same model used in recent studies (Ughini et al., 2022; Wu et al., 2023). To provide a more comprehensive insight into our results, we conduct visual comparisons among multiple examples. In summary, we present the primary contributions of our work:
  • We propose adversarial training on purification (AToP), a novel defense technique that effectively combines the strengths of adversarial training and adversarial purification.
  • We divide adversarial purification into two components, including perturbation destruction and fine-tuning purifier model, and investigate their impact on robustness.
@@ 2 Related work
Adversarial training is initially proposed by Goodfellow et al. (2015) to defend against adversarial attacks. While adversarial training has been demonstrated as an effective defense against attacks, it remains susceptible to unseen attacks (Stutz et al., 2020; Poursaeed et al., 2021; Laidlaw et al., 2021; Tack et al., 2022). Additionally, retraining the classifier model leads to substantial training expenses and could severely impair the standard accuracy of the model (Madry et al., 2018a; Tramèr et al., 2018; Zhang et al., 2019). In a study conducted by Raff et al. (2019), the impact of random transforms on the robustness of defense method was investigated. Their research revealed that the inclusion of a broader range of random transforms applied to images before adversarial training can significantly improve robustness. However, the utilization of random transforms also led to the loss of semantic information, which can severely affect the accuracy. Although our framework uses random transforms and AT strategy, the essential difference with these works is that we aim to train the purifier model to recover semantics from corrupted examples using AT, rather than the classifier model to learn the perturbation features using AT. Yang et al. (2019) also proposed a method that combines adversarial training and purification. Specifically, a purifier model was used prior to the classifier model, and AT was only applied to the classifier. Meanwhile, Ryu & Choi (2022) applied a similar pipeline but fine-tuned both the classifier model and the purifier model using AT. Consequently, the training process became time-intensive. In contrast, our method applies the adversarial loss to train the purifier model, which avoids the need for retraining the classifier model and mitigating the risk of overfitting specific attacks.
@@
Adversarial purification purify adversarial examples before classification, which has emerged as a promising defense method (Shi et al., 2021; Srinivasan et al., 2021). Currently, the most prevalent pipeline within AP involves the utilization of the generator model to recover purified examples from the adversarial examples. Some prior studies (Bakhti et al., 2019; Hwang et al., 2019) employed consistency losses to train the purifier model, which obtains similar results to AT methods. Nevertheless, these methods primarily capture the distribution of specific perturbations, which also limits their effectiveness in defending against unseen attacks. Other studies (Nie et al., 2022; Ughini et al., 2022; Wu et al., 2023) employed the pre-trained generator model as a purifier model. These methods have acquired the ability to recover information from noisy images through extensive exposure to clean examples, enabling them to provide an effective defense against even unseen attacks. However, there is a limitation of the existing pre-trained generator-based purifier model, where all parameters are fixed and cannot be further improved for known attacks (Dai et al., 2022). Considering these issues, we employ a pre-trained generator model as an initial purifier model, which can be subsequently fine-tuned using examples in a supervised manner, i.e., adversarial training on purifier model. Additionally, we use adversarial loss derived from classifier outputs rather than consistency loss derived from purifier outputs. In the experiment, our method achieves improved robustness accuracy without compromising generalization against unseen attacks.
@@ 3 Method
In this section, we introduce a novel defense technique called Adversarial Training on Purification (AToP), which comprises two components: random transform (RT) and fine-tuning (FT) the purifier model. Specifically, RT aims to destruct adversarial perturbations, irrespective of the type of attack, thereby achieving effective defense against even unseen attacks. On the other hand, FT aims to generate purified examples from corrupted inputs with ground truth, thus further enhancing robustness.
The entire classification system comprises three operations: transform denoted $t(\cdot)$, purifier denoted $g(\cdot)$, and classifier denoted $f(\cdot)$. For an input $x$, the pipeline of output $y$ can be formulated as
@@ 3.1 Destruct perturbation structure by random transforms
Randomized smoothing by adding Gaussian noise is a method that provides provable defenses (Li et al., 2019; Cohen et al., 2019) where Cohen et al. (2019) theoretically provides robustness guarantees for classifiers. In the latest work, Wu et al. (2023) enhance robustness by learning representations through a generator model, as shown in Eq. 3.
@@ 3.2 Fine-tuning the purifier model with adversarial loss
3:     Build adversarial examples $x^{\prime}$ with adversarial attacks:  $x^{\prime}\leftarrow x+\delta$
@@
Having described the random transforms, we proceed to provide a description of the purifier model and fine-tuning process. Specifically, based on pre-trained generator-based purifier model trained by the original loss function $L_{org}$, we have redesigned the loss function, incorporating an adversarial loss to fine-tune the purifier model:
@@
where $\lambda$ is held constant as hyperparameter. In the paper, we improve upon two recent studies, and the original loss function $L_{org}$ is denoted as $L_{df}$ in the GAN-based model (Ughini et al., 2022), and as $L_{mae}$ in the AE-based model (Wu et al., 2023). For instance, we utilize the GAN model called DeepFill as the purifier model $g$, in conjunction with $\rm{RT_{3}}$ and the classifier model $\rm{F}$. Given any input $x$, our defense process is defined as below:
@@
where $\theta_{t}$ contains three components: Gaussian standard $\sigma$, missing rate $r$, and mask number $N$. The parameter $\theta_{f}$ is derived from the pre-trained classifier model. To perform robust fine-tuning of the purifier model using adversarial training, we aim to optimize the model by freezing classifier parameter $\theta_{f}$ and updating purifier parameter $\theta_{g}$, where $\theta_{f}$ and $\theta_{g}$ are initialized by the pre-trained classifier model and the pre-trained generator model, respectively, as shown in Algorithm 1.
For pre-training GAN-based model (Ughini et al., 2022), we have $L_{\theta_{g}}=L_{df}\left(x,\theta_{g}\right)=\mathbb{E}[D(x)]-\mathbb{E}[D(g(t(x),\theta_{g}))]+\mathbb{E}[{\|x-g(t(x),\theta_{g})\|}_{\ell_{1}}]$. To further optimize $\theta_{g}$, we propose our AToP loss function and fine-tune the pre-trained GAN-based model, which comprise an original loss $L_{df}$, as well as an additional adversarial loss $L_{cls}$. For fine-tuning GAN-based model, we have
@@
$+\lambda\cdot\max_{\delta}CE\left\{y,\operatorname{F}[g(t(x^{\prime}),\theta_{g})]\right\},where\,x^{\prime}=x+\delta.$
Here, $L_{df}$ comprises the first three terms, where $D$ represents the discriminator, and $g$ is the generator-based purifier model, which plays similar roles as the generative adversarial network model. During training, the discriminator $D$ is responsible for distinguishing between real examples and the purified examples $g(t(x),\theta_{g})$. Simultaneously, the purifier model $g$ generates high-quality examples to deceive the discriminator $D$. The third term is consistency loss ${\|x-g(t(x),\theta_{g})\|}_{\ell_{1}}$, which encourages strong image restoration ability to the original examples $x$. Consequently, $L_{df}$ effectively ensures that the purified examples are not only high-quality but also closely resemble its original example. If RT is sufficiently strong to the extent that $t(x)$ is essentially Gaussian noise, $L_{df}$ loss can ensure the purified examples are generated from the same distribution of $x$ and are of high-quality. Conversely, if RT is minimal such as $t(x)=x$, $L_{df}$ is instrumental in guaranteeing that the purifier model can achieve an exact recovery of the original examples. Consequently, if RT is strong enough to destruct the adversarial perturbations, the model can guarantee to generate examples from the original distribution while avoiding the reconstruction of adversarial perturbations. Thus, the effectiveness remains unaffected even for unseen attacks.
During fine-tuning, $L_{cls}$ corresponds to the last term, which serves the purpose of further ensuring that our purified examples $\hat{x}$ do not contain sufficient adversarial perturbations that fool the classifier. On the other hand, if RT is very strong and purified examples $\hat{x}$ do not contain any adversarial perturbations, $L_{cls}$ can also ensure $\hat{x}$ will not be high-quality examples drawn from different classes. Intuitively, the adversarial perturbation $\delta$ is generated by maximizing classification loss, which can be used as the new examples to train our purifier model such that adversarial perturbations can be reconstructed up to the region that has correct classification. Differ from standard AT, our adversarial loss is not used to update classifier model but rather to optimize the purifier model. As compared to existing AP methods using adversarial examples (Bakhti et al., 2019; Hwang et al., 2019), our adversarial loss is essentially more robust than either consistency loss or generator-style loss. Finally, $\lambda>0$ represents a weight hyperparameter to balance the reconstruction error and classification error.

### 15 Mar 2024 v3

@@ Abstract
The deep neural networks are known to be vulnerable to well-designed adversarial attacks. The most successful defense technique based on adversarial training (AT) can achieve optimal robustness against particular attacks but cannot generalize well to unseen attacks. Another effective defense technique based on adversarial purification (AP) can enhance generalization but cannot achieve optimal robustness. Meanwhile, both methods share one common limitation on the degraded standard accuracy. To mitigate these issues, we propose a novel pipeline to acquire the robust purifier model, named Adversarial Training on Purification (AToP), which comprises two components: perturbation destruction by random transforms (RT) and purifier model fine-tuned (FT) by adversarial loss. RT is essential to avoid overlearning to known attacks, resulting in the robustness generalization to unseen attacks, and FT is essential for the improvement of robustness. To evaluate our method in an efficient and scalable way, we conduct extensive experiments on CIFAR-10, CIFAR-100, and ImageNette to demonstrate that our method achieves optimal robustness and exhibits generalization ability against unseen attacks.
@@ 1 Introduction
Deep neural networks (DNNs) have been shown to be susceptible to adversarial examples (Szegedy et al., 2014; Deng et al., 2020), which are generated by adding small, human-imperceptible perturbations to natural images, but completely change the prediction results to DNNs, leading to disastrous implications (Goodfellow et al., 2015). Since then, numerous methods have been proposed to defend against adversarial examples. Notably, adversarial training (AT) has gained significant attention as a robust learning algorithm against adversarial attacks (Goodfellow et al., 2015; Madry et al., 2018a). AT has proven to be effective in achieving state-of-the-art robustness against known attacks, albeit with a tendency to overfit the specific adversarial examples seen during training. Consequently, within the framework of AT, the persistent conundrum revolves around striking a balance between standard accuracy and robustness (Zhang et al., 2019), a dilemma that has been a subject of enduring concern. Furthermore, the robustness of AT defense against multiple or even unforeseen attacks still remains challenging (Poursaeed et al., 2021; Laidlaw et al., 2021; Tack et al., 2022).
More recently, the concept of adversarial purification (AP) has emerged, aiming to eliminate perturbations from potentially attacked images before inputting them into the classifier model (Yang et al., 2019; Shi et al., 2021; Nie et al., 2022). In contrast to AT methods, AP operates as a pre-processing step that can defend against unseen attacks without retraining the classifier model. However, AP exhibits lower robustness against known attacks compared to AT methods. Meanwhile, both AT and AP have a common limitation on the degraded standard accuracy (Tramer et al., 2020; Chen et al., 2022), which restricts their applicability in real-world applications. As summarized in Table 1, the present landscape of defense methods is confronted with a formidable challenge: How can we enhance robustness against known attacks while maintaining generalization to unseen attacks and preserving standard accuracy on clean examples?
@@
To tackle these challenges with the framework of AT and AP, we propose a novel defense technique that conceptually separates the AP method into two components: perturbation destruction by random transforms and fine-tuning the purifier model via supervised AT called Adversarial Training on Purification (AToP), as shown in Figure 1. Specifically, we develop three types of random transforms and conduct a comprehensive investigation into their effectiveness in augmenting robustness and generalization against attacks. The purifier model is designed to reconstruct clean examples from corrupted inputs, which can also ensure accurate classification outputs regardless of whether the inputs are corrupted adversarial examples or corrupted clean examples. Furthermore, more importantly, we introduce the adversarial loss calculated from the output of the classifier model to fine-tune the purifier model in a supervised manner while keeping the classifier model fixed. Differing from the traditional AP methods, our method can learn a robust purifier model to generate high-quality purified examples and avoid generating examples to incorrect classes.
We empirically evaluate the performance of our method by comparing with the latest AT and AP methods across various attacks (i.e., FGSM (Goodfellow et al., 2015), PGD (Madry et al., 2018b), CW (Carlini & Wagner, 2017), AutoAttack (Croce & Hein, 2020), and StAdv (Xiao et al., 2018)) and conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette employing multiple classifier models and purifier models. The experimental results demonstrate that our method achieves state-of-the-art results and exhibits robust generalization against unseen attacks. Furthermore, our method significantly improves the performance of the purifier model in robust classification, surpassing the performance of the same model used in recent studies (Ughini et al., 2022; Wu et al., 2023). To provide a more comprehensive insight into our results, we conduct visual comparisons among multiple examples. In summary, the contributions of this work are as follows.
  • We propose adversarial training on purification (AToP), a novel defense technique that effectively combines the strengths of AT and AP to acquire the robust purifier model.
  • We separate the adversarial purification process into two components, including perturbation destruction and fine-tuning purifier model, and investigate their impacts on robustness.
@@ 2 Related work
Adversarial training is initially proposed by Goodfellow et al. (2015) to defend against adversarial attacks. While adversarial training has been demonstrated as an effective defense against attacks, it remains susceptible to unseen attacks (Stutz et al., 2020; Poursaeed et al., 2021; Laidlaw et al., 2021; Tack et al., 2022). Additionally, retraining the classifier model leads to substantial training expenses and could severely impair the standard accuracy of the model (Madry et al., 2018a; Tramèr et al., 2018; Zhang et al., 2019). In a study conducted by Raff et al. (2019), the impact of random transforms on the robustness of defense method was investigated. Their research revealed that the inclusion of a broader range of random transforms applied to images before adversarial training can significantly improve robustness. However, the utilization of random transforms also led to the loss of semantic information, which can severely affect the accuracy. Although our framework uses random transforms and AT strategy, the essential difference with these works is that we aim to train a robust purifier model to recover semantics from corrupted examples using AT, rather than training the classifier model to learn the perturbation features using AT. Yang et al. (2019) also proposed a method that combines adversarial training and purification. Specifically, a purifier model was used prior to the classifier model, and AT was only applied to the classifier. Meanwhile, Ryu & Choi (2022) applied a similar pipeline but fine-tuned both the classifier model and the purifier model using AT. Consequently, the training process became time-intensive. In contrast, our method applies the adversarial loss to train the purifier model, which avoids the need for retraining the classifier model and mitigating the risk of overfitting specific attacks.
@@
Adversarial purification aims to purify adversarial examples before classification, which has emerged as a promising defense method (Shi et al., 2021; Srinivasan et al., 2021). Currently, the most prevalent pipeline within AP involves the utilization of a generative model to recover purified examples from adversarial examples. Some prior studies (Bakhti et al., 2019; Hwang et al., 2019) employed consistency losses to train the purifier model, which obtains similar results to AT methods. Nevertheless, these methods primarily capture the distribution of specific perturbations, which also limits their effectiveness in defending against unseen attacks. Other studies (Nie et al., 2022; Ughini et al., 2022; Wu et al., 2023) employed the pre-trained generative model as a purifier model. These methods have acquired the ability to recover information from noisy images through extensive exposure to clean examples, enabling them to provide an effective defense against even unseen attacks. However, there is a limitation of the existing pre-trained generator-based purifier model, where all parameters are fixed and cannot be further improved for known attacks (Dai et al., 2022). To address these issues, we employ a pre-trained generative model as an initial purifier model, which can be subsequently fine-tuned using examples in a supervised manner, i.e., adversarial training on purifier model. Additionally, we introduce an adversarial loss derived from classifier outputs rather than consistency loss derived from purifier outputs. As shown in the experiments, our method achieves improved robustnessness without compromising generalization against unseen attacks.
@@ 3 Method
In this section, we introduce a novel defense pipeline to acquire the robust purifier model, named Adversarial Training on Purification (AToP), which comprises two components: random transform (RT) and fine-tuning (FT) the purifier model. Specifically, RT can significantly destruct adversarial perturbations, irrespective of the type of attack, thereby achieving effective defense against even unseen attacks. On the other hand, FT can generate purified examples from corrupted inputs with true label information, thus further enhancing robustness.
The entire classification system comprises three processes: transform denoted $t(\cdot)$, purifier denoted $g(\cdot)$, and classifier denoted $f(\cdot)$. Given an input $x$, the pipeline of output $y$ can be formulated as
@@ 3.1 Destruct perturbation structure by random transforms
Randomized smoothing by adding Gaussian noise is a method that provides provable defenses (Li et al., 2019; Cohen et al., 2019) where Cohen et al. (2019) theoretically provides robustness guarantees for classifiers. In the latest work, Wu et al. (2023) enhance robustness by learning representations through a generative model, as shown in Eq. 3.
@@ 3.2 Fine-tuning the purifier model with adversarial loss
3:     Build adversarial examples $x^{\prime}$ with perturbations $\delta$:  $x^{\prime}\leftarrow x+\delta$
@@
Having described the random transforms, we proceed to provide a description of the purifier model and fine-tuning process. Specifically, based on the pre-trained generator-based purifier model trained by the original loss function $L_{org}$, we propose AToP loss function $L_{\theta_{g}}$, incorporating an adversarial loss to fine-tune the purifier model,
@@
where $\lambda$ is held constant as hyperparameter. In the paper, we improve upon two recent studies, and the original loss function $L_{org}$ is denoted as $L_{df}$ in the GAN-based model (Ughini et al., 2022) and as $L_{mae}$ in the AE-based model (Wu et al., 2023). For instance, we utilize the GAN-based model called DeepFill as the purifier model $g$, in conjunction with $\rm{RT_{3}}$ and the classifier model $\rm{F}$. Given any input $x$, our defense process is defined as below:
@@
where $\theta_{t}$ contains three components: Gaussian standard $\sigma$, missing rate $r$, and mask number $N$. The parameter $\theta_{f}$ is derived from the pre-trained classifier model. To perform robust fine-tuning of the purifier model using adversarial training, we aim to optimize the model by freezing classifier parameter $\theta_{f}$ and updating purifier parameter $\theta_{g}$, where $\theta_{f}$ and $\theta_{g}$ are initialized by the pre-trained classifier model and the pre-trained generative model, respectively, as shown in Algorithm 1.
For pre-training GAN-based model (Ughini et al., 2022), we have $L_{org}=L_{df}\left(x,\theta_{g}\right)=\mathbb{E}[D(x)]-\mathbb{E}[D(g(t(x),\theta_{g}))]+\mathbb{E}[{\|x-g(t(x),\theta_{g})\|}_{\ell_{1}}]$. To further optimize $\theta_{g}$, we propose AToP loss function and fine-tune the pre-trained GAN-based model, which comprises an original loss $L_{df}$, as well as an additional adversarial loss $L_{cls}$. Given an adversarial example $x^{\prime}$ with perturbation $\delta$, we have
@@
$+\lambda\cdot\max_{\delta}\,CE\left\{y,\operatorname{F}[g(t(x^{\prime}),\theta_{g})]\right\},\text{where}\,x^{\prime}=x+\delta.$
Here, $L_{df}$ comprises the first three terms, where $D$ represents the discriminator, and $g$ is the generator-based purifier model, which plays similar roles as the generative adversarial network model. During training, the discriminator $D$ is responsible for distinguishing between real examples and the purified examples $g(t(x),\theta_{g})$. Simultaneously, the purifier model $g$ generates high-quality examples to deceive the discriminator $D$. The third term is consistency loss ${\|x-g(t(x),\theta_{g})\|}_{\ell_{1}}$, which encourages strong image restoration ability to the original examples $x$. Consequently, $L_{df}$ effectively ensures that the purified examples are not only high-quality but also closely resemble their original example. If RT is sufficiently strong to the extent that $t(x)$ is essentially Gaussian noise, $L_{df}$ loss can ensure the purified examples are generated from the same distribution of $x$ and are of high-quality. Conversely, if RT is minimal such as $t(x)=x$, $L_{df}$ is instrumental in guaranteeing that the purifier model can achieve an exact recovery of the original examples. Consequently, if RT is strong enough to destruct the adversarial perturbations, the model can guarantee to generate examples from the original distribution while avoiding the reconstruction of adversarial perturbations. Thus, the effectiveness remains unaffected even for unseen attacks.
During fine-tuning, $L_{cls}$ corresponds to the last term, which serves the purpose of further ensuring that our purified examples $\hat{x}$ do not contain sufficient adversarial perturbations that fool the classifier. On the other hand, if RT is very strong and purified examples $\hat{x}$ do not contain any adversarial perturbations, $L_{cls}$ can also ensure $\hat{x}$ will not be high-quality examples drawn from different classes. Intuitively, the adversarial perturbations $\delta$ are generated by maximizing classification loss, which can be used as the new examples to train our purifier model such that adversarial perturbations can be reconstructed up to the degree that remains correct classification. Unlike standard AT, our adversarial loss is not used to update the classifier model but to optimize the purifier model. As compared to existing AP methods using adversarial examples (Bakhti et al., 2019; Hwang et al., 2019), our adversarial loss is essentially more robust than either consistency loss or generator-style loss. Finally, $\lambda>0$ represents a weight hyperparameter to balance the generative error and classification error.